OpenStudy (anonymous):

What is an equation for the line that passes through points (–1, –4) and (1, 4)? Write the equation in slope-intercept form

OpenStudy (lgbasallote):

well..there is such an equation known as.. y-y1 = [(y2-y1)/(x2-x1)](x-x1) x1 = -1 y1 = -4 x2 = 1 y2 = 4 substitute the values then transform to slope-intercept... which of the steps did you find confusing??

OpenStudy (anonymous):

no idea what you just said

OpenStudy (anonymous):

\[\LARGE y-y_1=\frac{y_2-y_1}{x_2-x_1}\cdot (x-x_1)\] you have: \[\LARGE (\underbrace{-1}_{x_1}\;\;\;,\;\;\;\underbrace{-4}_{y_1})\] \[\LARGE (\underbrace{1}_{x_2}\;\;\;,\;\;\;\underbrace{4}_{y_2})\] now substitute...

OpenStudy (anonymous):

whattttt?

OpenStudy (anonymous):

@xitskaylee23 you should be able to do it now, I don't see anything confusing.

OpenStudy (anonymous):

unless if you don't even have the idea about these problems,and just asked to see if someone will do it for you.... !

OpenStudy (anonymous):

yup, im totally lost. im cyber schooled and its hard trying to learn geometry on your own

OpenStudy (anonymous):

well then I suggest you to google these problems... there you'll find tons of them (examples) http://www.analyzemath.com/line/Tutorials.html